# Unleashing Diffusion Transformers for Visual Correspondence by Modulating Massive Activations

### v1

@@ 6 Conclusion
In this paper, we identify and characterize massive activations in Diffusion Transformers (DiTs). We observe that these massive activations are concentrated in a few fixed dimensions across all image patch tokens and carry limited local information. We further demonstrate that the built-in AdaLN mechanism in DiTs effectively suppresses massive activations while enhancing feature semantics and discriminability. Building on these insights, we propose a training-free framework, Diffusion Transformer Feature (DiTF), which extracts semantically discriminative features from DiTs via channel-wise modulation through AdaLN. Extensive experiments on visual correspondence tasks validate the robustness and effectiveness of our approach.
@@ Appendix A Diffusion Transformer Architecture
We strictly follow the architecture used in DiT [36]. We provide an illustration of the DiT architecture as shown in Figure 9.
@@ B.1 Massive Activations in DiTs
To further illustrate the emergence of massive activations in DiTs, we provide additional visualization results in Figure 13. We show the LayerNorm-normalized activation magnitudes of original features from SD2-1 and various DiTs. While SD2-1 exhibits smooth activations, DiTs consistently show spikes concentrated in a few fixed dimensions across all patch tokens, revealing a fundamental difference that contributes to their degraded performance.
@@ B.2 Massive Activations Dimensions Align with Residual Scaling Factor
In this section, we provide additional visualizations to examine the dimensional alignment between massive activations and the residual scaling factor $\alpha_{k}$ from the AdaLN layer. As illustrated in Figure 10, massive activations consistently co-occur with large values of $\alpha_{k}$ in the same dimensions across all DiTs.
@@ B.3 Channel-wise Modulation with AdaLN
To comprehensively demonstrate the impact of the built-in AdaLN in DiTs, we provide additional visualizations comparing pre-AdaLN and post-AdaLN features across various models, including SD3-5 (Figure 11), Pixart-Alpha (Figure 14), SD3 (Figure 15), and Flux (Figure 16). These results consistently show that AdaLN accurately localizes and normalizes massive activations, while enhancing feature semantics and discrimination through effective channel-wise modulation.
@@ Appendix C Further Implementation Details
Configurations of different DiTs. We employ the pre-trained Diffusion Transformers (DiTs) as a feature extractor for semantic correspondence. Formally, we decompose the universal feature extraction process in DiTs into two stages: (1) extracting the original feature $z_{t}^{k}$ from the DiT block $\mathcal{A}_{k}$, and (2) modulating it via adaptive channel-wise scaling and shifting using the AdaLN layer. For Pixart-alpha [21], SD3 [22], and SD3-5 [22], we first extract the pre-AdaLN feature $z_{t}^{(k,2)}$ and then activate it as follows.
@@
As some of the Flux [23] model’s blocks contain only one group of AdaLN-zero layer, we extract pre-AdaLN feature $z_{t}^{(k,1)}$ and then activate it as follows.
@@
The configurations of different DiTs can be found in Table 7, where the total time step T is 1000. We set the input image size as 960x960 for DiTs and 840x840 for the model DINOv2.
Investigation on block index $k$ and timestep $t$. The layer index $k$ and the timestep $t$ are critical hyperparameters that influence the quality of the extracted features from Diffusion Transformers (DiTs). Previous studies [18, 14] have conducted thorough investigations to identify the optimal layer index $k$ and timestep $t$ for Stable Diffusion. To explore the effects of varying $k$ and $t$ in DiTs, we conducted grid search experiments to identify the optimal hyperparameters. The results are presented in Figure 12. The figure reveals that features extracted from the middle layer achieve optimal performance across different DiTs. Furthermore, feature extraction in DiTs is robust to the timestep for semantic correspondence tasks, as a wide range of $t$ achieves excellent performance.
@@
SIFT [26] None 40.2 68.0 79.3 26.8 55.4 72.1 54.6 81.5 86.9
[…]
SuperPoint [61] 36.4 72.7 82.6 22.1 56.1 68.2 51.9 90.8 98.1
D2-Net [62] 16.7 61.0 75.9 3.7 38.0 56.6 30.2 84.9 95.8
DISK [63] 40.2 70.6 81.5 23.2 51.4 67.9 58.5 91.2 96.2
ContextDesc [64] 40.9 73.0 82.2 29.6 60.7 72.5 53.1 86.2 92.7
R2D2 [65] 40.0 74.4 84.3 26.4 60.4 73.9 54.6 89.6 95.4
@@
CAPS [66] Weak 44.8 76.3 85.2 35.7 62.9 74.3 54.6 90.8 96.9
[…]
OpenCLIP [67] 33.3 67.2 78.0 18.6 45.0 59.6 49.2 91.2 97.7
DIFT [18] 45.6 73.9 83.1 30.4 56.8 69.3 61.9 92.3 98.1
[…]
@@
MAST YT-VOS [68] 65.5 63.3 67.6
SFC [69] 71.2 68.3 74.0
InstDis [68] ImageNet [70] 66.4 63.9 68.9
MoCo [71] 65.9 63.4 68.4
SimCLR [72] 66.9 64.4 69.4
BYOL [73] w/o labels 66.5 64.0 69.0
SimSiam [74] 67.2 64.8 68.8
DINO [32] 71.4 67.9 74.9
[…]
DIFT [18] 70.0 67.4 72.5
[…]
@@ Appendix D Geometric Correspondence
Datasets. Following [18], we evaluate our model on the HPatches benchmark [76], which comprises 116 sequences: 57 with illumination changes and 59 with viewpoint variations. Adopting the approach from CAPS [66], we detect up to 1,000 key points per image and apply cv2.findHomography() to estimate homography using mutual nearest neighbor matches.
Metric. We adopt the corner correctness metric for evaluation where we compute the average error between the four estimated corners of one image and the ground-truth corners with a threshold $\epsilon$ pixels, following [66, 18],.
[…]
@@ Appendix E Temporal Correspondence
Datasets. We conduct experiments on the challenge video dataset: DAVIS-2017 video instance segmentation benchmark [77], following [18].
Metric. Following [18, 78], we adopt the region-based similarity $\mathcal{J}$ and contour-based accuracy $\mathcal{F}$ as the performance metric where we segment the nearest neighbors between the consecutive video frames based on the representation similarity.
Results. The temporal correspondence results can be found in Table 9. From the results, it can be observed that our model exhibits a superior capability to extract video frame features for temporal correspondence. Specifically, our model achieves 72.2% on the dataset DAVIS, which outperforms the previous state-of-the-art DIFT by 2.2%, demonstrating the superior effectiveness of our model.
@@ Appendix F Qualitative Results on AP-10K
We show the qualitative comparison of our Diffusion Transformer model DiTF with both Stable Diffusion (SD) and SD+DINO [14] in AP-10K intra-species (Figure 17), cross-species ( Figure 18), and cross-family ( Figure 19) subset.
Appendix G Limitations and Future Work
Lightweight Fine-Tuning for Enhanced Representation. In this work, we propose a training-free framework that leverages the built-in AdaLN in DiTs to suppress massive activations and enhance feature semantics and discrimination. While our approach demonstrates strong performance without additional training, further research could explore lightweight fine-tuning strategies to better adapt DiTs for representation learning and fully unlock their potential as feature extractors.
Understanding Massive Activations in Generation Tasks. In this work, we identify and characterize massive activations in Diffusion Transformers (DiTs) from a representation learning perspective, aiming to mitigate their undesirable impact on discriminative feature extraction. Our focus is orthogonal to the generative aspect of DiTs. We believe that further exploration of massive activations from the viewpoint of visual generation could yield valuable insights and potentially enhance generative performance.

### v3

@@ 6 Conclusion
In this paper, we identify and characterize massive activations in Diffusion Transformers (DiTs). We observe that these activations consistently emerge in a few fixed dimensions across all image patch tokens and carry limited local information. We further demonstrate that the built-in AdaLN mechanism in DiTs effectively suppresses massive activations while enhancing feature semantics and discriminability. Building on these insights, we propose a training-free AdaLN-based framework, Diffusion Transformer Feature (DiTF), which extracts semantically discriminative features from DiTs through channel-wise modulation using AdaLN. Extensive experiments on visual correspondence tasks validate the robustness and effectiveness of our approach.
7 Limitations and Future Work.
In this work, we primarily focus on understanding and mitigating massive activations in Diffusion Transformers (DiTs) from the perspective of representation learning. However, the emergence and potential roles of massive activations in the generative process of DiTs remain underexplored. We believe that investigating these activations from a generative viewpoint could provide deeper insights into their underlying mechanisms and potentially contribute to improving generative performance.
Acknowledgements
The paper is supported in part by the National Natural Science Foundation of China (No. U21B2013, 62325109), and in part by the Shanghai ’The Belt and Road’ Young Scholar Exchange Grant (24510742000). Mehrtash Harandi is supported by the Australian Research Council (ARC) Discovery Program DP250100262.
@@ Appendix A Diffusion Transformer Architecture
We strictly follow the architecture used in DiT [peebles2023scalable]. We provide an illustration of the DiT architecture as shown in Figure˜9.
@@ B.1 Massive Activations in DiTs
To further illustrate the emergence of massive activations in DiTs, we provide additional visualization results in Figure˜13. We show the LayerNorm-normalized activation magnitudes of original features from SD2-1 and various DiTs. While SD2-1 exhibits smooth activations, DiTs consistently show spikes concentrated in a few fixed dimensions across all patch tokens, revealing a fundamental difference that contributes to their degraded performance.
@@ B.2 Massive Activations Dimensions Align with Residual Scaling Factor
In this section, we provide additional visualizations to examine the dimensional alignment between massive activations and the residual scaling factor $\alpha_{k}$ from the AdaLN layer. As illustrated in Figure˜10, massive activations consistently co-occur with large values of $\alpha_{k}$ in the same dimensions across all DiTs.
@@ B.3 Channel-wise Modulation with AdaLN
To comprehensively demonstrate the impact of the built-in AdaLN in DiTs, we provide additional visualizations comparing pre-AdaLN and post-AdaLN features across various models, including SD3-5 (Figure˜11), Pixart-Alpha (Figure˜14), SD3 (Figure˜15), and Flux (Figure˜16). These results consistently show that AdaLN accurately localizes and normalizes massive activations, while enhancing feature semantics and discrimination through effective channel-wise modulation.
@@ Appendix C Further Implementation Details
Configurations of different DiTs. We employ the pre-trained Diffusion Transformers (DiTs) as a feature extractor for semantic correspondence. Formally, we decompose the universal feature extraction process in DiTs into two stages: (1) extracting the original feature $z_{t}^{k}$ from the DiT block $\mathcal{A}_{k}$, and (2) modulating it via adaptive channel-wise scaling and shifting using the AdaLN layer. For Pixart-alpha [chen2023pixart], SD3 [esser2024scaling], and SD3-5 [esser2024scaling], we first extract the pre-AdaLN feature $z_{t}^{(k,2)}$ and then activate it as follows.
@@
As some of the Flux [Flux] model’s blocks contain only one group of AdaLN-zero layer, we extract pre-AdaLN feature $z_{t}^{(k,1)}$ and then activate it as follows.
@@
The configurations of different DiTs can be found in Table˜7, where the total time step T is 1000. We set the input image size as 960x960 for DiTs and 840x840 for the model DINOv2.
Investigation on block index $k$ and timestep $t$. The layer index $k$ and the timestep $t$ are critical hyperparameters that influence the quality of the extracted features from Diffusion Transformers (DiTs). Previous studies [tang2023emergent, zhang2024tale] have conducted thorough investigations to identify the optimal layer index $k$ and timestep $t$ for Stable Diffusion. To explore the effects of varying $k$ and $t$ in DiTs, we conducted grid search experiments to identify the optimal hyperparameters. The results are presented in Figure˜12. The figure reveals that features extracted from the middle layer achieve optimal performance across different DiTs. Furthermore, feature extraction in DiTs is robust to the timestep for semantic correspondence tasks, as a wide range of $t$ achieves excellent performance.
@@
SIFT [lowe2004distinctive] None 40.2 68.0 79.3 26.8 55.4 72.1 54.6 81.5 86.9
LF-Net [ono2018lf] Strong 34.4 62.2 73.7 16.8 43.9 60.7 53.5 81.9 87.7
SuperPoint [detone2018superpoint] 36.4 72.7 82.6 22.1 56.1 68.2 51.9 90.8 98.1
D2-Net [dusmanu2019d2] 16.7 61.0 75.9 3.7 38.0 56.6 30.2 84.9 95.8
DISK [tyszkiewicz2020disk] 40.2 70.6 81.5 23.2 51.4 67.9 58.5 91.2 96.2
ContextDesc [luo2019contextdesc] 40.9 73.0 82.2 29.6 60.7 72.5 53.1 86.2 92.7
R2D2 [revaud2019r2d2] 40.0 74.4 84.3 26.4 60.4 73.9 54.6 89.6 95.4
@@
CAPS [wang2020learning] Weak 44.8 76.3 85.2 35.7 62.9 74.3 54.6 90.8 96.9
DINO [oquab2023dinov2] None 38.9 70.0 81.7 21.4 50.7 67.1 57.7 90.8 97.3
OpenCLIP [ilharco_gabriel_2021_5143773] 33.3 67.2 78.0 18.6 45.0 59.6 49.2 91.2 97.7
DIFT [tang2023emergent] 45.6 73.9 83.1 30.4 56.8 69.3 61.9 92.3 98.1
$\text{DiTF}_{\texttt{flux}}\text{(ours)}$} 41.9 70.7 79.5 22.0 50.8 63.4 62.5 91.3 96.2
@@
MAST YT-VOS [xu2018youtube] 65.5 63.3 67.6
SFC [hu2022semantic] 71.2 68.3 74.0
InstDis [xu2018youtube] ImageNet [deng2009imagenet] 66.4 63.9 68.9
MoCo [he2020momentum] 65.9 63.4 68.4
SimCLR [chen2020simple] 66.9 64.4 69.4
BYOL [grill2020bootstrap] w/o labels 66.5 64.0 69.0
SimSiam [chen2021exploring] 67.2 64.8 68.8
DINO [caron2021emerging] 71.4 67.9 74.9
OpenCLIP [ilharco_gabriel_2021_5143773] LAION [schuhmann2022laion] 62.5 60.6 64.4
DIFT [tang2023emergent] 70.0 67.4 72.5
$\text{DiTF}_{\texttt{flux}}\text{(ours)}$} 72.2 69.2 75.1
@@ Appendix D Geometric Correspondence
Datasets. Following [tang2023emergent], we evaluate our model on the HPatches benchmark [balntas2017hpatches], which comprises 116 sequences: 57 with illumination changes and 59 with viewpoint variations. Adopting the approach from CAPS [wang2020learning], we detect up to 1,000 key points per image and apply cv2.findHomography() to estimate homography using mutual nearest neighbor matches.
Metric. We adopt the corner correctness metric for evaluation where we compute the average error between the four estimated corners of one image and the ground-truth corners with a threshold $\epsilon$ pixels, following [wang2020learning, tang2023emergent],.
Results. To comprehensively evaluate our models, we conducted geometric correspondence experiments on the HPatches benchmark [balntas2017hpatches], as detailed in Table˜8. From the results, it can be observed that our model enables robust feature extraction for image pairs and draws precise geometric correspondence. Specifically, our model $\text{DiTF}_{\texttt{flux}}\text{(ours)}$} achieve comparable performance 41.9% compared to the state-of-the-art model DIFT (SD-based). These results show that Diffusion Transformers can be employed as an effective feature extractor for geometric correspondence.
@@ Appendix E Temporal Correspondence
Datasets. We conduct experiments on the challenge video dataset: DAVIS-2017 video instance segmentation benchmark [pont20172017], following [tang2023emergent].
Metric. Following [tang2023emergent, li2019joint], we adopt the region-based similarity $\mathcal{J}$ and contour-based accuracy $\mathcal{F}$ as the performance metric where we segment the nearest neighbors between the consecutive video frames based on the representation similarity.
Results. The temporal correspondence results can be found in Table˜9. From the results, it can be observed that our model exhibits a superior capability to extract video frame features for temporal correspondence. Specifically, our model achieves 72.2% on the dataset DAVIS, which outperforms the previous state-of-the-art DIFT by 2.2%, demonstrating the superior effectiveness of our model.
@@ Appendix F Qualitative Results on AP-10K
We show the qualitative comparison of our Diffusion Transformer model DiTF with both Stable Diffusion (SD) and SD+DINO [zhang2024tale] in AP-10K intra-species (Figure˜17), cross-species ( Figure˜18), and cross-family ( Figure˜19) subset.